### v1

@@ Abstract.
Conformal prediction (CP) is an Uncertainty Representation technique that delivers finite-sample calibrated prediction regions for any underlying Machine Learning model, yet its status as an Uncertainty Quantification (UQ) tool has remained conceptually opaque. We adopt a category-theoretic approach to CP—framing it as a morphism, embedded in a commuting diagram, of two newly-defined categories—that brings us three joys. First, we show that—under minimal assumptions—CP is intrinsically a UQ mechanism, that is, its UQ capabilities are a structural feature of the method. Second, we demonstrate that CP bridges (and perhaps subsumes) the Bayesian, frequentist, and imprecise probabilistic approaches to predictive statistical reasoning. Finally, we show that a conformal prediction region (CPR) is the image of a covariant functor. This observation is relevant to AI privacy: It implies that privacy noise added locally does not break coverage.
@@ 1. Introduction
Conformal Prediction (CP, introduced formally in Section 2.1) is a popular ML prediction tool. Suppose we are given an $n$-sized exchangeable training data $y^{n}=(y_{1},\ldots,y_{n})$ where each observation $y_{i}$ belongs to the state space $Y$, a confidence level $\alpha\in[0,1]$, and a non-conformity score $\psi$ (i.e. a function that measures “how different” a new observation is with respect to the ones in the training set). CP, then, returns a prediction set, called the Conformal Prediction Region (CPR), that contains the “true value” $y_{n+1}$ with high probability $1-\alpha$ (Vovk et al.,, 2005; Shafer and Vovk,, 2008; Vovk,, 2013; Angelopoulos et al.,, 2024). [^2] For extensions, and more modern references, see e.g. (Caprio et al., 2025b, ; Javanmardi et al.,, 2024; Gibbs et al.,, 2023; Barber et al.,, 2023; Papadopoulos,, 2008). In Caprio et al., 2025a (, Section 5.1), the authors show how CP can be seen as a correspondence (i.e. a set-valued function), thus highlighting a connection between ML and functional analysis (FA, Aliprantis and Border, (2006)). Other such connections have been explored in the literature, see e.g. Korolev, (2022).
@@
  • CP is not only an uncertainty representation, but also an uncertainty quantification tool: The scholars who improperly referred to CP as an Uncertainty Quantification (UQ) methodology, were right all along.
  • CP bridges Bayesian, frequentist, and imprecise probabilistic approaches to predictive statistical reasoning, thus formalizing an intuition first put forth by Martin, (2022).
  • Seeing the Conformal Prediction Region as the image of a covariant functor directly benefits AI privacy (Fang et al.,, 2025): (i) Any locally added privacy noise leaves coverage guarantees intact, and (ii) In federated learning (Liu et al.,, 2024), each agent need only share a summary object (the credal set $\mathcal{M}$ that we introduce in the following paragraphs) instead of raw data.
@@
Not only is the CP correspondence a morphism of these categories, but it is also part of a diagram that commutes. As we shall see (in particular in Example 1 and in the discussion thereafter), this discovery implies the intrinsic UQ capabilities of CP.
@@ 2.2. Imprecise Probabilities
The conjugate lower probability $\underline{P}$ of $\overline{P}$ satisfies $\underline{P}(A)=1-\overline{P}(A^{c})$, for all $A\in\Sigma_{Y}$, thus knowing one immediately gives us the other. In a statistical and ML framework, these concepts allow us to crystallize model ambiguity.
@@
The consonance assumption is needed to ensure that $\mathsf{CRED}(y^{n},\psi)\in\mathscr{C}$. In the remainder of the paper, we sometimes write $\pi(y,y^{n},\psi)\equiv\pi_{\psi}(y,y^{n})$ for notational convenience (we drop the $n+1$ subscript for $y_{n+1}$), and to explicitly acknowledge that the conformal transducer $\pi$ depends on the choice of the non-conformity measure $\psi$.
We also consider a slight modification of $\mathsf{CRED}$, that will considerably simplify our proving endeavors in Sections 3, 4, and 5; We denote it by $\mathsf{CRED}_{\text{cl}}(y^{n},\psi)\coloneqq\{P\in\Delta_{Y}:P(A)\leq\overline{\Pi}(A)\text{, }\forall A\in\Sigma_{Y}\text{ closed}\}$. As we can see, $\mathsf{CRED}_{\text{cl}}(y^{n},\psi)\supseteq\mathcal{M}(\overline{\Pi})$, because it imposes less constraints on the $P$’s: They only have to be setwise dominated on closed sets. The upper probability $\overline{\Pi}$ and its conjugate lower probability $\underline{\Pi}$ (the only characteristics of the credal sets that we use) stay the same for both $\mathsf{CRED}$ and $\mathsf{CRED}_{\text{cl}}$: This justifies our simplifying choice. To avoid notation complications, in the remainder of this work we do not distinguish between $\mathsf{CRED}(y^{n},\psi)$ and $\mathsf{CRED}_{\text{cl}}(y^{n},\psi)$.
@@ 3. Conformal Prediction as an Upper Hemicontinuous Morphism
(iii) $\mathsf{IHDR}$. For each $A\in\Sigma_{Y}$ define $f_{A}(\alpha,\mathcal{M})\coloneqq\underline{P}_{\!\mathcal{M}}(A)-(1-\alpha)$. Weak continuity of $P\mapsto P(A)$ and compactness of $\mathcal{M}$ imply that $\mathcal{M}\mapsto\underline{P}_{\!\mathcal{M}}(A)$ is upper semicontinuous in the Vietoris topology (iv), while $\alpha\mapsto 1-\alpha$ is continuous. Hence the set
@@
is closed; Moreover $(\alpha,\mathcal{M})\mapsto H_{A}(\alpha,\mathcal{M})$ is closed-valued by (i). In turn, $\operatorname{Gr}(\mathsf{IHDR})=\bigl{\{}(\alpha,\mathcal{M},y):y\in Y\setminus\bigcap_{A}H_{A}(\alpha,\mathcal{M})\bigr{\}}^{c}=\bigcap_{A}(H_{A}\times Y)$ is closed.
@@
Hence, $\psi_{\theta}\in\mathscr{F}$. Notice that consonance, i.e. $\sup_{y\in Y}\pi(y,y^{n})=1$, does not automatically hold in this example. To enforce it, it is enough to consider the normalized conformal transducer $\tilde{\pi}(y,y^{n})\coloneqq\frac{\pi(y,y^{n})}{\sup_{y\in Y}\pi(y,y^{n})}$ (Cella and Martin,, 2021, Section 7), which is still upper semicontinuous.
@@
The commuting diagram, however, reveals a hidden layer: Computing a CP region is equivalent to (i) forming the credal set $\mathcal{M}(\overline{\Pi})$ and (ii) extracting its IHDR. Because $\mathcal{M}(\overline{\Pi})$ lives in the space of probability measures, existing metrics—e.g. those in Abellán et al., (2006); Hüllermeier and Waegeman, (2021); Javanmardi et al., (2023); Hofman et al., (2024)—endow it with a cardinal scale that is able to quantify (different types of) uncertainty, in particular, reducible (epistemic) and irreducible (aleatoric) (Javanmardi et al.,, 2024, 2025; Cabezas et al.,, 2025; Stutts et al.,, 2024; Hanselle et al.,, 2025). This will be “marginal uncertainty”, [^8] That is, averaged over the possible realization $y^{n}$. thus reflecting the marginal guarantee of the Conformal Prediction Region (see equation (2)). We conjecture that the analyst-controlled choices—the significance level $\alpha$, which fixes the tolerated error, and the non-conformity score $\psi\in{\mathscr{F}}$, which embodies modeling assumptions—are primarily associated with epistemic uncertainty. On the other hand, aleatoric uncertainty stems from the intrinsic variability of the exchangeable process that we study, encoded in $\mathfrak{P}$.
@@
Remark 1 (Minimality of Hypotheses (i)-(iv)).
@@
We prove that each hypothesis is indispensable for the simultaneous upper hemicontinuity (u.h.c.) of the three correspondences $\kappa$, $\mathsf{CRED}$, and $\mathsf{IHDR}$. For every clause, we remove the hypothesis in question while retaining the others, and we show that at least one correspondence fails to be upper hemicontinuous.
  • Drop (i) (compactness of $Y$). Let $Y=\mathbb{R}$ with the Euclidean topology (non-compact), $n=1$, and keep $\psi(y^{1},y)=-y$ (continuous, so (ii) holds). Fix $\alpha=\tfrac{3}{4}\notin\{0,\tfrac{1}{2},1\}$ (satisfies the no-tie rule). Writing $y^{1}=(x)$,
    $\kappa(\alpha,y^{1},\psi)=\{y\in\mathbb{R}:y\geq x\}\;=\;[x,\infty).$
    Take the sequence $x_{k}\nearrow x_{0}=1$. For the open set $V\coloneqq(1,\infty)\subset\kappa(\alpha,x_{0},\psi)=[1,\infty)$, no neighborhood of $(\alpha,x_{0},\psi)$ keeps $\kappa(\alpha,x_{k},\psi)\subseteq V$ (the left end-points $x_{k}$ lie outside $V$). Thus $\kappa$ is not u.h.c. at $(\alpha,x_{0},\psi)$.
  • Drop (ii) (joint continuity of $\psi$). Keep $Y=[0,1]$ (compact) and $n=1$. Define the pathological non-conformity score
    $\psi(y^{1},y)=\begin{cases}1&\text{if }y\in\mathbb{Q}\\ 0&\text{otherwise}\end{cases},$
    which is permutation-invariant but nowhere continuous. Take $\alpha=\tfrac{1}{2}$, $y^{1}=(0)$. Then, $\pi(y,y^{1})=1$ for rational $y$’s and $\pi(y,y^{1})=\tfrac{1}{2}$ for irrational $y$’s, so $\kappa(\alpha,y^{1},\psi)=\mathbb{Q}\cap[0,1]$, a dense set that is not closed. Because nearby bags $\bigl{(}y^{1}+\varepsilon\bigr{)}$ shuffle rationals and irrationals, the graph of $\kappa$ fails to be closed at $(\alpha,y^{1},\psi)$, hence $\kappa$ is not u.h.c.
@@
  • Drop (iii) (no-tie rule). Work again with the continuous $\psi_{\theta}$ of Example 1, $n\geq 1$, and choose $\alpha=\tfrac{1}{n+1}$, which is in $S_{n+1}$. Let $y^{n}$ be such that all non-conformity scores tie (e.g. $y=y_{1}=\cdots=y_{n}$), so $\pi(y,y^{n})=\tfrac{1}{n+1}$, for every $y$. Then, $\kappa(\alpha,y^{n},\psi_{\theta})=\emptyset$, whereas for the sequence $\tfrac{1}{n+1}-\tfrac{1}{k}\eqqcolon\alpha_{k}\downarrow\alpha$ we have $\kappa(\alpha_{k},y^{n},\psi_{\theta})=Y$. Thus $\kappa$ is not u.h.c. at $\alpha$ unless the no-tie exclusion is imposed.
  • Drop (iv) (weak topology on $\Delta_{Y}$). Still with $Y=[0,1]$, equip $\Delta_{Y}$ with the total-variation norm. For any fixed closed Borel set $A$ the map $P\mapsto P(A)$ is not continuous in total variation (take Dirac masses sliding along $A$). Consequently, the constraint $P(A)\leq\overline{\Pi}(A)$ fails to define a closed set in $Y^{n}\times\mathscr{F}\times\Delta_{Y}$, so the graph of $\mathsf{CRED}$ is no longer closed and upper hemicontinuity breaks down.
    Drop (iv) (Vietoris topology on $\mathscr{C}$). Keep the weak topology on $\Delta_{Y}$, but put on $\mathscr{C}\subset 2^{\Delta_{Y}}$ the Hausdorff-metric topology instead of Vietoris. Take $Y=\{0,1\}$ so that $\Delta_{Y}=[0,1]$ via $p\mapsto P(\{1\})=p$, and consider the event $A=\{1\}$. Define the sequence of credal sets
    $\mathcal{M}_{k}=[0,1-\tfrac{1}{k}],\qquad\mathcal{M}_{\infty}=\{1\}.$
    The Hausdorff distance $d_{H}(\mathcal{M}_{k},\mathcal{M}_{\infty})=1/k\to 0$, so $\mathcal{M}_{k}\to\mathcal{M}_{\infty}$ in the Hausdorff topology, but
[…]
    Thus $\mathcal{M}\mapsto\underline{P}_{\!\mathcal{M}}(A)$ is not upper semicontinuous, so $\mathsf{IHDR}$ fails to be u.h.c.
Of course, consonance is needed to apply Caprio et al., 2025a (, Proposition 5). Therefore, (i)-(iv), together with consonance, are collectively minimal.

### v2

@@ Abstract.
Conformal prediction (CP) is an Uncertainty Representation technique that delivers finite-sample calibrated prediction regions for any underlying Machine Learning model. Its status as an Uncertainty Quantification (UQ) tool, though, has remained conceptually opaque: While Conformal Prediction Regions (CPRs) give an ordinal representation of uncertainty (larger regions typically indicate higher uncertainty), they lack the capability to cardinally quantify it (twice as large regions do not imply twice the uncertainty). We adopt a category-theoretic approach to CP—framing it as a morphism, embedded in a commuting diagram, of two newly-defined categories—that brings us three joys. First, we show that—under minimal assumptions—CP is intrinsically a UQ mechanism, that is, its cardinal UQ capabilities are a structural feature of the method. Second, we demonstrate that CP bridges (and perhaps subsumes) the Bayesian, frequentist, and imprecise probabilistic approaches to predictive statistical reasoning. Finally, we show that a CPR is the image of a covariant functor. This observation is relevant to AI privacy: It implies that privacy noise added locally does not break the global coverage guarantee.
@@ 1. Introduction
Conformal Prediction (CP, introduced formally in Section 2.1) is a popular ML prediction tool. Suppose we are given an $n$-sized exchangeable training data $y^{n}=(y_{1},\ldots,y_{n})$ where each observation $y_{i}$ belongs to the state space $Y$, a confidence level $\alpha\in[0,1]$, and a non-conformity score $\psi$ (i.e. a function that measures “how different” a new observation is with respect to the ones in the training set). CP, then, returns a prediction set, called the Conformal Prediction Region (CPR), that contains the “true value” $y_{n+1}$ with high probability $1-\alpha$ (Vovk et al.,, 2005; Shafer and Vovk,, 2008; Vovk,, 2013; Angelopoulos et al.,, 2024). [^2] For extensions, and more modern references, see e.g. (Caprio et al., 2025b, ; Javanmardi et al.,, 2024; Gibbs et al.,, 2023; Barber et al.,, 2023; Papadopoulos,, 2008). In Caprio et al., 2025a (, Section 5.1), the authors show how CP can be seen as a correspondence (i.e. a set-valued function, sometimes referred to as relation (Perrone,, 2024)), thus highlighting a connection between ML and functional analysis (FA, Aliprantis and Border, (2006)). Other such connections have been explored in the literature, see e.g. Korolev, (2022).
@@
  • Quantified uncertainty. CP is not only an uncertainty (ordinal) representation, but also an uncertainty (cardinal) quantification tool: The scholars who improperly referred to CP as an Uncertainty Quantification (UQ) methodology, were right all along.
    • As a byproduct, the user will now be able to see (i) how the choice of non-conformity measure $\psi$ controls the size of the prediction region, and (ii) how much it affects (different types of) uncertainty in the analysis.
  • Conceptual Unification. CP bridges Bayesian, frequentist, and imprecise probabilistic approaches to predictive statistical reasoning, thus formalizing an intuition first put forth by Martin, (2022). This is a conceptual contribution that shows the central (and possibly unifying) role that CP plays in prediction theory.
  • Privacy by Design. Seeing the Conformal Prediction Region as the image of a covariant functor directly benefits AI privacy (Fang et al.,, 2025): (i) Any locally added privacy noise leaves coverage guarantees intact, and (ii) In federated learning (Liu et al.,, 2024), each agent need only share a summary object (the credal set $\mathcal{M}$, introduced in the following paragraphs) instead of raw data.
@@
Not only is the CP correspondence a morphism of these categories, but it is also part of a diagram that commutes. As we shall see (in particular in Example 1 and in the discussion thereafter), this discovery implies the intrinsic (cardinal) UQ capabilities of CP.
@@ 2.2. Imprecise Probabilities
The conjugate lower probability $\underline{P}_{\mathcal{M}}\equiv\underline{P}$ of $\overline{P}$ satisfies $\underline{P}(A)=1-\overline{P}(A^{c})$, for all $A\in\Sigma_{Y}$, thus knowing one immediately gives us the other. In a statistical and ML framework, these concepts allow us to crystallize model ambiguity.
@@
The consonance assumption is needed to ensure that $\mathsf{CRED}(y^{n},\psi)\in\mathscr{C}$; The elements of $\mathsf{CRED}(y^{n},\psi)$ are characterized in Martin, (2025, Theorem 1). In the remainder of the paper, we sometimes write $\pi(y,y^{n},\psi)\equiv\pi_{\psi}(y,y^{n})$ for notational convenience (we drop the $n+1$ subscript for $y_{n+1}$), and to explicitly acknowledge that the conformal transducer $\pi$ depends on the choice of the non-conformity measure $\psi$.
We also consider a slight modification of $\mathsf{CRED}$, that will significantly simplify our proving endeavors in Sections 3, 4, and 5; We denote it by $\mathsf{CRED}_{\text{cl}}(y^{n},\psi)\coloneqq\{P\in\Delta_{Y}:P(A)\leq\overline{\Pi}(A)\text{, }\forall A\in\Sigma_{Y}\text{ closed}\}$. As we can see, $\mathsf{CRED}_{\text{cl}}(y^{n},\psi)\supseteq\mathcal{M}(\overline{\Pi})$, because it imposes less constraints on the $P$’s: They only have to be setwise dominated on closed sets. The upper probability $\overline{\Pi}$ and its conjugate lower probability $\underline{\Pi}$ (the only characteristics of the credal sets that we use) stay the same for both $\mathsf{CRED}$ and $\mathsf{CRED}_{\text{cl}}$: This justifies our simplifying choice. To avoid notation complications, in the remainder of this work we do not distinguish between $\mathsf{CRED}(y^{n},\psi)$ and $\mathsf{CRED}_{\text{cl}}(y^{n},\psi)$.
@@ 3. Conformal Prediction as an Upper Hemicontinuous Morphism
(iii) $\mathsf{IHDR}$. For each $A\in\Sigma_{Y}$ define $f_{A}(\alpha,\mathcal{M})\coloneqq\underline{P}_{\!\mathcal{M}}(A)-(1-\alpha)$. Weak continuity of $P\mapsto P(A)$ and compactness of $\mathcal{M}$ imply that $\mathcal{M}\mapsto\underline{P}_{\!\mathcal{M}}(A)$ is lower semicontinuous in the Vietoris topology (iv), while $\alpha\mapsto 1-\alpha$ is continuous. Hence the set
@@
is closed; Moreover $(\alpha,\mathcal{M})\mapsto H_{A}(\alpha,\mathcal{M})$ is closed-valued by (i). In turn, $\operatorname{Gr}(\mathsf{IHDR})=\cap_{A\in\Sigma_{Y}}(([0,1]\times\mathscr{C})\setminus H_{A})\times Y\cup(H_{A}\times A)$ is closed.
@@
Hence, $\psi_{\theta}\in\mathscr{F}$. Notice that consonance, i.e. $\sup_{y\in Y}\pi(y,y^{n})=1$, does not automatically hold in this example. To enforce it, it is enough to consider the normalized conformal transducer $\tilde{\pi}(y,y^{n})\coloneqq\frac{\pi(y,y^{n})}{\sup_{y\in Y}\pi(y,y^{n})}$ (Cella and Martin,, 2021, Section 7), which is still upper semicontinuous because the denominator is a strictly positive constant for fixed $y^{n}$.
@@
The commuting diagram, however, reveals a hidden layer: Computing a CP region is equivalent to (i) forming the credal set $\mathcal{M}(\overline{\Pi})$ and (ii) extracting its IHDR. Because $\mathcal{M}(\overline{\Pi})$ lives in the space of probability measures, existing metrics—e.g. those in Abellán et al., (2006); Hüllermeier and Waegeman, (2021); Javanmardi et al., (2023); Hofman et al., (2024); Chau et al., (2025)—endow it with a cardinal scale that is able to quantify (different types of) uncertainty, in particular, reducible (epistemic) and irreducible (aleatoric) (Javanmardi et al.,, 2024, 2025; Cabezas et al.,, 2025; Stutts et al.,, 2024; Hanselle et al.,, 2025). This will be “marginal uncertainty”, [^8] That is, averaged over the possible realization $y^{n}$. thus reflecting the marginal guarantee of the Conformal Prediction Region (see equation (2)). We conjecture that the analyst-controlled choices—the significance level $\alpha$, which fixes the tolerated error, and the non-conformity score $\psi\in{\mathscr{F}}$, which embodies modeling assumptions—are primarily associated with epistemic uncertainty. On the other hand, aleatoric uncertainty stems from the intrinsic variability of the exchangeable process that we study, encoded in $\mathfrak{P}$.
@@
Remark 1 (Minimality of Hypotheses (i)(iv)).
@@
We prove that each hypothesis is indispensable for the simultaneous upper hemicontinuity (u.h.c.) of the three correspondences $\kappa$, $\mathsf{CRED}$, and $\mathsf{IHDR}$. For each clause, we remove exactly one hypothesis and exhibit a single parameter value at which at least one of the three maps fails to be u.h.c. [^9] We use u.h.c. for both “upper hemicontinuity” and “upper hemicontinuous”. Throughout, all remaining hypotheses (plus consonance) continue to hold.
  1. (i)
    Dropping (i) (compactness of $Y$). Let $Y=\mathbb{R}$ with its usual, non-compact topology, take $n=1$, and define $\psi(y^{1},y)\equiv-y$ (jointly continuous, so (ii) stays true). Fix the closed set $A\!\coloneqq(-\infty,0]$. For $y^{1}=(x)$ a short calculation gives
    $\overline{\Pi}_{x}(A)=\begin{cases}1,&x\leq 0,\\[2.0pt] \tfrac{1}{2},&x>0.\end{cases}$
    Hence $x\mapsto\overline{\Pi}_{x}(A)$ has a downward jump at $x=0$, so the envelope is not u.s.c. Consequently $\mathsf{CRED}$ (and therefore $\mathsf{IHDR}$) is not u.h.c. Compactness of $Y$ is thus indispensable.
  2. (ii)
    Dropping (ii) (joint continuity of $\psi$). Let $Y=[0,1]$ (compact Hausdorff) and $n=1$. Define the permutation-invariant but nowhere continuous score
    $\psi(y^{1},y)=\begin{cases}1,&y\in\mathbb{Q},\\ 0,&y\notin\mathbb{Q}.\end{cases}$
    Fix $\alpha=\tfrac{3}{4}$ and $y^{1}=(0)$. Then $\pi(y,y^{1})=1\text{ if }y\notin\mathbb{Q},$ and $\pi(y,y^{1})=\tfrac{1}{2}\text{ if }y\in\mathbb{Q}.$ Hence $\kappa(\alpha,y^{1},\psi)=[0,1]\cap\mathbb{Q}^{c}$, which is dense but not closed. Because nearby samples $y^{1}=(x)$ with $x\downarrow 0$ randomly interchange rationals and irrationals, the graph of $\kappa$ fails to be closed at $(\alpha,y^{1},\psi)$, so $\kappa$ is not u.h.c. Joint continuity of $\psi$ is therefore necessary.
  3. (iii)
@@
    Dropping (iii) (no-tie rule). Work with any continuous $\psi\in\mathscr{F}$ (e.g. the neural $\psi_{\theta}$ of Example 1) and fix $n\geq 1$. Choose $\alpha=1\in S_{n+1}$ (which violates the no-tie exclusion). If all $n$ observed points coincide, then every candidate $y$ ties with the sample, giving $\pi(y,y^{n})=1$ for all $y$. Because $\kappa$ uses the strict inequality $\pi>\alpha$, we obtain $\kappa(1,y^{n},\psi)=\emptyset$. For the sequence $1-\tfrac{1}{k}\eqqcolon\alpha_{k}\downarrow 1$ we have $\kappa(\alpha_{k},y^{n},\psi)=Y$ for every $k$, so $\kappa$ is not u.h.c. at $\alpha=1$. The no-tie parameter restriction is thus essential.
  4. (iv)
    Dropping (iv) (topologies on $\Delta_{Y}$ and $\mathscr{C}$).
    1. (a)
      Replace the weak topology on $\Delta_{Y}$ by the total-variation (TV) norm. Although individual evaluations $P\mapsto P(A)$ remain 1-Lipschitz in TV, $\Delta_{Y}$ is no longer compact, so the Closed Graph Theorem (Lemma 3) breaks down. A concrete example: Take $Y=[0,1]$, fix any $(y^{n},\psi)$, and consider the sequence of Dirac measures $P_{k}\coloneqq\delta_{1/k}\in\mathsf{CRED}(y^{n},\psi)$. The graph-points $\bigl{(}(y^{n},\psi),P_{k}\bigr{)}$ have no TV-convergent subsequence in $(Y^{n}\times\mathscr{F})\times\Delta_{Y}$, so $\mathsf{CRED}$ fails to be u.h.c. Compactness of $\Delta_{Y}$ in the weak topology is therefore indispensable.
    2. (b)
      Keep the weak topology on $\Delta_{Y}$ but equip $\mathscr{C}$ with the Hausdorff metric instead of the Vietoris topology. Let $Y=\{0,1\}$ so that $\Delta_{Y}=[0,1]$ via $p\mapsto P(\{1\})$, and write $A\coloneqq\{1\}$. Define $\mathcal{M}_{k}\coloneqq[0,1-\tfrac{1}{k}]$ and $\mathcal{M}_{\infty}\coloneqq\{1\}$. Then $d_{H}(\mathcal{M}_{k},\mathcal{M}_{\infty})=\tfrac{1}{k}\to 0$, yet $\underline{P}_{\!\mathcal{M}_{k}}(A)=0\to 0<1=\underline{P}_{\!\mathcal{M}_{\infty}}(A),$ so $\underline{P}_{\mathcal{M}}(A)$ is not lower semicontinuous, and $\mathsf{IHDR}$ is not u.h.c. The Vietoris topology on $\mathscr{C}$ is therefore indispensable.
Of course, consonance is needed to apply Caprio et al., 2025a (, Proposition 5). In short, removing any one of (i)-(iv) (while keeping the rest, plus consonance) makes at least one of $\kappa$, $\mathsf{CRED}$, or $\mathsf{IHDR}$ non-u.h.c.; The hypotheses are collectively minimal. $\triangle$